# Bayesian decomposition using Besov priors

### v1

@@ 4.2 One dimensional deconvolution: Hierarchical decomposition method
First, we test the hierarchical decomposition method with two Besov priors on this problem. The hierarchical posterior and the sampling scheme are given in A. We use the Haar wavelet for $\psi_{g}$ and the DB(8) wavelet for $\psi_{h}$, with prior smoothness parameters set to $s_{g}=1$ and $s_{h}=3$, respectively. In the Gibbs sampler, we set $a_{1}=a_{2}=2$ and $b_{1}=b_{2}=10^{-3}$. We draw 1 000 00010000001\,000\,0001 000 000 samples and use the first 10 0001000010\,00010 000 samples as burn-in, and thin the remaining samples by a factor of $200$, resulting in $4950$ samples for posterior analysis.
@@ 4.3 Image deblurring
In this section, we apply our hierarchical decomposition method proposed in Section 3 on a 2D deconvolution problem. We consider a separable Gaussian convolution kernel, i.e., the convolution is performed through a 1D Gaussian convolution defined in Section 4.1 along both the horizontal and vertical directions. The ground truth is the sum of a piecewise constant component and a smooth component. The resolution of $\mathbf{f}_{\text{true}}$ is $n=64^{2}=$4096$$, and the intensity range is $[0,2]$. To generate the degraded data, we set $\sigma_{\text{ker}}=1.0$ and the noise level $0.02$. Figure 10 shows both the ground truth and the degraded image.
@@
In our method, we use the DB(8) wavelet with $s_{h}=3$ in the Besov prior for $\psi_{h}$. The parameters in the hyperpriors are set as $\alpha_{1}=1$, $\alpha_{2}=2$, $\beta_{1}=10^{-5}$ and $\beta_{2}=10^{-2}$. We run Algorithm 1 to generate 200 000200000200\,000200 000 samples, from which we discard the first 10 0001000010\,00010 000 as burn-in and thin the remaining samples by a factor of 30, providing $6334$ samples used for analysis.
@@ Appendix B Anisotropic uncertainty in Besov priors
This covariance matrix is not a multiple of the identity unless $s+d/2-d/p=0$ and the wavelet transform is orthonormal, i.e., bi-orthogonal wavelets or non-trivial Besov weights will lead to anisotropic covariance.

### v2

@@ 4.2 One dimensional deconvolution: Hierarchical decomposition method
First, we test the hierarchical decomposition method with two Besov priors on this problem. The hierarchical posterior and the sampling scheme are given in A. We use the Haar wavelet for $\psi_{g}$ and the DB(8) wavelet for $\psi_{h}$, with prior smoothness parameters set to $s_{g}=1$ and $s_{h}=3$, respectively. In the Gibbs sampler, we set $a_{1}=a_{2}=2$ and $b_{1}=b_{2}=10^{-3}$. We draw 1 000 0001\,000\,000 samples and use the first 10 00010\,000 samples as burn-in, and thin the remaining samples by a factor of $200$, resulting in $4950$ samples for posterior analysis.
@@ 4.3 Image deblurring
In this section, we apply our hierarchical decomposition method proposed in Section 3 on a 2D deconvolution problem. We consider a separable Gaussian convolution kernel, i.e., the convolution is performed through a 1D Gaussian convolution defined in Section 4.1 along both the horizontal and vertical directions. The ground truth is the sum of a piecewise constant component and a smooth component. The resolution of $\mathbf{f}_{\text{true}}$ is $n=64^{2}=$4096$\$, and the intensity range is $[0,2]$. To generate the degraded data, we set $\sigma_{\text{ker}}=1.0$ and the noise level $0.02$. Figure 10 shows both the ground truth and the degraded image.
@@
In our method, we use the DB(8) wavelet with $s_{h}=3$ in the Besov prior for $\psi_{h}$. The parameters in the hyperpriors are set as $\alpha_{1}=1$, $\alpha_{2}=2$, $\beta_{1}=10^{-5}$ and $\beta_{2}=10^{-2}$. We run Algorithm 1 to generate 200 000200\,000 samples, from which we discard the first 10 00010\,000 as burn-in and thin the remaining samples by a factor of 30, providing $6334$ samples used for analysis.
@@ Appendix B Anisotropic uncertainty in Besov priors
This covariance matrix is not a multiple of the identity unless s+d/2d/p=0s+d/2-d/p=0 and the wavelet transform is orthonormal, i.e., bi-orthogonal wavelets or non-trivial Besov weights will lead to anisotropic covariance.